OpenStudy (anonymous):

May you help me, please :) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

how far did you get with this?

OpenStudy (anonymous):

I got 14 rolls. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

same here 1 wall: 7*7 = 49 2 walls: 49*2 = 98 so 2 walls = 98 sq ft ------------------------------------------------------- 1 roll: 10*0.7 = 7 sq ft x rolls: 7*x = 7x sq ft 7x = 98 x = 98/7 x = 14 so you need 14 rolls

OpenStudy (anonymous):

That's what I got, yay :)